OpenStudy (anonymous):

An auto race consists of 8 laps. A driver completes the first 3 laps at an average speed of 185 miles/hour and the remaining laps at an average speed of 200 miles/hour. If d represents the length of one lap, which expression represents the time in hours, t, that it takes the driver to complete the race?

OpenStudy (nurali):

welcome to openstudy.

OpenStudy (anonymous):

thanks!

OpenStudy (nurali):

option please.

OpenStudy (nurali):

reply

OpenStudy (anonymous):

8d/192.5 = t 5d/185 + 3d/200 = t 8d/385 = t 3d/185 + 5d/200 = t

OpenStudy (nurali):

d = r*t If the race is 8 laps and the total miles is 500, that means that each lap represents 62.5 miles. If the driver does the first 3 laps at 185 miles per hour, we use d = r*t. Each lap is symbolized by d = 62.5 miles. Using the formula, we now have the equation below to find to do the first 3 laps. 3*d/(185) Do the same thing for the next part: 5*d/(200) Add them together and you get t. The answer is D.

OpenStudy (anonymous):

THANK YOU SO MUCH!! Your awesome :)

OpenStudy (nurali):

Anytime.